### v2

# High-pressure characterization of Ag3AuTe2: Implications for strain-induced band tuning

Abstract
Recent band structure calculations have suggested the potential for band tuning in a chiral semiconductor, Ag3AuTe2, to zero upon application of negative strain. In this study, we report on the synthesis of polycrystalline Ag3AuTe2 and investigate its transport, optical properties, and pressure compatibility. Transport measurements reveal the semiconducting behavior of Ag3AuTe2 with high resistivity and an activation energy $E_{a}$ of 0.2 eV. The optical band gap determined by diffuse reflectance measurements is about three times wider than the experimental $E_{a}$. Despite the difference, both experimental gaps fall within the range of predicted band gaps by our first-principles DFT calculations employing the PBE and mBJ methods. Furthermore, our DFT simulations predict a progressive narrowing of the band gap under compressive strain, with a full closure expected at a strain of -4% relative to the lattice parameter. To evaluate the feasibility of gap tunability at such substantial strain, the high-pressure behavior of Ag3AuTe2 was investigated by in situ high-pressure X-ray diffraction up to 47 GPa. Mechanical compression beyond 4% resulted in a pressure-induced structural transformation, indicating the possibilities of substantial gap modulation under extreme compression conditions.
I Introduction
Metal chalcogenides have emerged as a remarkable class of materials, exhibiting unique electronic properties that make them suitable for exploring topological phenomena. Petzite or Ag3AuTe2 is a mineral, discovered by Petz about two centuries agoFaizan et al. (2017). It belongs to the uytenbogaardtite mineral group along with Ag3AuSe2 and Ag3AuS2. While Ag3AuSe2 and Ag3AuTe2 adopt the cubic $I4_{1}32$ space group, Ag3AuS2 takes on the tetragonal $P4_{1}22$ structure.Barton et al. (1978) Frueh et al.Frueh (1959) elucidated the crystal structure of natural Ag3AuTe2 by single crystal X-ray diffraction (XRD).
High-temperature powder XRD experiment revealed phase transitions in Ag3AuTe2 at 210 and 319 C, with an intermediate phase eluding identification.Cabri (1965); Frueh (1959) These transitions were corroborated by differential thermal analysis.Tavernier et al. (1967); Smit et al. (1970) The initial synthesis of Ag3AuTe2 was reported by Thompson,Thompson (1948) providing limited details on temperature or purity, except for conducting reactions in evacuated silica tubes. Young et al.Young et al. (2000) investigated the thermoelectric properties of synthetic Ag3AuTe2  noting low thermal conductivity and large electrical resistivity.
Recent computational predictions have pointed to a narrow and direct gap at $\Gamma$ Faizan et al. (2016); Young et al. (2000); Sánchez-Martínez et al. (2019), opening possibilities for non-trivial topological behavior such as band crossing and inversion via band gap tuning. Sánchez-Martínez et al.Sánchez-Martínez et al. (2019) explored potential applications of Ag3AuSe2 and Ag3AuTe2 in dark matter detection, highlighting their tunable band gap through strain. Subsequent band closure of Ag3AuTe2 was predicted with a 4% reduction in the lattice parameter,Sánchez-Martínez et al. (2019) attainable through approaches such as chemical substitution and application of hydrostatic pressure. Our preceding computational examination of the band gap tunability of Ag3AuSe2 demonstrated a narrowing of the gap under tensile strainWon et al. (2022), aligning with findings by Sánchez-Martínez et al.Sánchez-Martínez et al. (2019) However, such heightened strain on a material is subject to deformation, decomposition, or fracture. Therefore, a thorough assessment of the mechanical compressibility of Ag3AuTe2 is necessary to evaluate the feasibility of band engineering by achieving -4% strain.
Previous studies on Ag3AuTe2Faizan et al. (2016); Young et al. (2000); Sánchez-Martínez et al. (2019) highlight a small direct gap with potential band closure under compressive strain, yet experimental determination of the band gap and compressibility remains scant. Here, we report the synthesis, transport, optical properties and mechanical compressibility of Ag3AuTe2 and explore the DFT-simulated band structure at varying lattice constants.
II Methods
Powders of Ag (99.9% metals basis), Au (99.98% metals basis), and Te (99.8% metals basis) were loaded and sealed in an evacuated silica ampule in a 3.05:1:1.9 ratio. Subsequently, the ampule underwent annealing in a muffle furnace at 770 C for a day, followed by rapid cooling to 200 C and a cool-down to room temperature over three days. This procedure yielded a brittle ingot exhibiting a metallic luster and covered partially with gold flakes on the surface. High-resolution powder X-ray diffraction (PXRD) data were collected using the beamline 11-BM at the Advanced Photon Source (APS), Argonne National Laboratory (ANL).Lee et al. (2008) Data were collected at 288 K, using a wavelength of 0.45897 Å and the capillary geometry. Rietveld refinement was performed utilizing GSAS-II.Toby and Von Dreele (2013)
High-pressure XRD measurements were performed at the 16-BMD-HPCAT beamline of the APS. Finely ground sample was loaded with a few grains of ruby and a Ne pressure-transmitting medium in a symmetric diamond anvil cell (DAC) with 300-$\mu$m culet diamonds and a Re gasket. The gaskets were indented to 60 $\mu$m, and a circular hole of 130 $\mu$m diameter was laser drilled to create the sample chamber. The Ne gas was loaded at the GSECARS facility at APS. The calibration of the sample center was carried out with a CeO2 standard, and the high-pressure data were collected up to 47 GPa with an incident beam of wavelength 0.41326 Å. Pressure at the sample region was measured and monitored by the shift of ruby fluorescence linesH.K.Mao and Bell (1986). The XRD images were collected with a Pilatus detector and integrated using Dioptas software Prescher and Prakapenka (2015).
Electrical resistivity measurements were performed by the 4-point contact probe method using the Quantum Design Physical Property Measurement System (PPMS). The Ag3AuTe2 ingot was cut and polished into a rod with dimensions 3.97 × 1.04 × 0.93 mm3 to facilitate transport measurements. Gold leads were attached to the sample with a two-part silver epoxy for the measurements.
Diffuse reflectance spectra of non-diluted Ag3AuTe2 powder were acquired using a Varian Cary 5G UV-Vis-NIR spectrophotometer in the spectral range of 200-3000 nm. A Spectralon pellet, sourced from Harrick Scientific Products, served as the reference standard.
The density-functional theory simulations were performed employing the Vienna Ab initio Simulation Package (VASP) codeKresse and Hafner (1993, 1994) with Perdew–Burke–Ernzerhof (PBE)Perdew et al. (1996) projector-augmented wave (PAW) pseudopotentials.Blöchl (1994); Hobbs et al. (2000) We used the cutoff energy of 500 eV, and $\Gamma$ centered $8\times 8\times 8$ grid in the self-consistent field calculation. Calculations with modified Becke-Johnson (mBJ) exchange-correlation functions Becke and Johnson (2006); Tran and Blaha (2009) were performed with default parameters $\alpha=-0.012$, $\beta=1.023$ bohr1/2, $e=0.500$ Tran and Blaha (2009). Experimental lattice constants in ambient pressure and compressive pressure were considered in the simulations. The high-symmetry path in the band structure calculations is defined as $\Gamma(0.0,0.00.0)$, $H(0.5,-0.5,0.5)$, $N(0.0,0.0,0.5)$, $P(0.25,0.25,0.25)$.
III Results and Discussion
III.1 Growth and structure
Figure 1: The crystal structure of Ag3AuTe2 is shown in (a) with the cubic unit cell indicated by the black box. The $4_{1}$ screw operation is well-demonstrated in (b) by the Ag chain spiraling into the plane.
Figure 2: Rietveld-refined synchrotron PXRD pattern of Ag3AuTe2. The contributions from minor phases are shown in the figure and quantified in the text.
The room temperature crystal structure of petzite ($a=10.38$, $Z=8$) is analogous to fischesserite, Ag3AuSe2, as described in detail in our previous work.Won et al. (2022) The crystal structure of Ag3AuTe2 features a chiral $4_{1}$ Ag chain down the $\langle 100\rangle$ axis, as illustrated in Figure 1. Our Ag3AuTe2 sample was characterized by PXRD and analysed using Rietveld method. As shown in Figure 2, Bragg peaks were well-indexed by the cubic petzite structure and some minor impurities (wR = 2.84%). Phase fractions of Ag3AuTe2, Ag2Te, Ag doped Au2Te and Au were refined to 0.926, 0.035, 0.034, and 0.005 by weight, respectively.
An equilibrated 3-component system can accommodate up to three phases in an isotherm, which denotes an disequilibrium in our final product. The Au-Ag2Te phase diagramCabri (1965) and previous studies on the thermal properties of Ag3AuTe2Tavernier et al. (1967); Smit et al. (1970); Cabri (1965); Frueh (1959) identify petzite as a line compound undergoing two phase transformations below its melting point. Despite our efforts to synthesize Ag3AuTe2 via a slow cooling method from the melt, incorporating adjustments in cooling rates and elemental ratios, all attempts resulted in a mixture of Ag3AuTe2, Ag2Te, Au2Te, and Au, contradicting the reported phase diagram. There are certain ambiguities in the phase diagram proposed by Cabri et al. Cabri (1965), such as the dashed line above the melting point of Ag3AuTe2 labeled with a question mark. In addition, it is unclear whether the ”liquid” above the melting point is a uniform phase or molten impurities. Furthermore, the Ag3AuTe2 phase behavior near 200 C is ambiguous, with an intermediate phase mentioned but not indexed. This prompts questions about potential phase separation or deviation from the nominal 312 stoichiometry, leading to phase separation around the transition temperature. Consequently, the reformation of the lowest form of Ag3AuTe2 upon cooling may be challenging due to the difficulty in reaching an equilibrium at such low temperatures.
Figure 3: (Top) in situ high-pressure XRD patterns of Ag3AuTe2 collected at various pressures ($\lambda=0.41326$). Le Bail fit for the orthorhombic ($Pnma$) structure at 47 GPa is displayed. (Bottom) Volume per formula unit of refined cells for Ag3AuTe2 plotted as a function of applied pressure. The multi-phase region around 16 GPa does not provide well-converged values for lattice parameters of the individual phases. The error bars for the volume are smaller than the symbol size.
Figure 4: Rietveld-refined lattice parameters and the corresponding $\sigma$ of compressed Ag3AuTe2 are plotted against the applied pressure. Nearly -4% of strain is achieved at 10 GPa.
III.2 High-Pressure X-ray Diffraction
The high-pressure behavior of Ag3AuTe2 was investigated through in situ high-pressure XRD experiments. Figure 3 depicts the XRD patterns collected at various representative pressures, reaching up to 47 GPa. The cubic structure of Ag3AuTe2 at ambient pressure remained stable up to 13 GPa, demonstrated by the rightward shift of the most intense (110) peak at 8.5. Beyond 13 GPa, this prominent peak drops in intensity, indicating a pressure-induced phase change. Notably, a number of peaks merge around 10 and new peaks emerge near 17.5. Rietveld analysis confirmed the retention of the ambient cubic phase up to 13 GPa and the resulting refined lattice constant $a$ showed a gradual decrease with increasing pressure, reaching a maximum strain exceeding -4% at 13 GPa, as shown Figure 4. Even though the patterns collected between 13 and 20 GPa show a mixture of both low- and high-pressure phases of Ag3AuTe2, the significant changes observed in the diffraction peaks around 16 GPa imply structural transformation. Analysis commenced using the 47 GPa pattern for the high-pressure phase corresponds mostly to the high-pressure phase.
We considered possible structural models from the pattern indexing and refinements using GSAS-II, LHPM Rietica, and QualXToby and Von Dreele (2013); Hunter (2000); Altomare et al. (2015) software programs as well as previously published papers for related compoundsBerri et al. (2013). Here, we consider an Al3CePt2-type hexagonal cell, $Pnma$ structure of CePd3In2,Nesterenko et al. (2004) and two high-pressure polymorphs of Ag2Te: $C2/m$ from Zhu, et al.Zhu et al. (2016) and $Pnma$ from Zhang, et al.Zhang et al. (2015) Attempts to fit the observed diffraction patterns with the above candidate structures favored the Ag2Te-type structure, replicating most observed peaks at 47 GPa, as shown in Figure 3. The resulting cell parameters were a = 13.452(2) Å, b = 3.531(8) Å, c = 4.188(1) Å. Fitting the patterns above 20 GPa with a similar cell resulted in a volume collapse of approximately 6% around the phase transition, adding weight to the likeliness that Ag3AuTe2 adopts the Pnma-type symmetry at high pressure. Ab initio enthalpy calculations on the proposed structural models would further shed light on their stability, especially since the number of observed diffraction peaks is small.
The bulk modulus of the low-pressure (cubic) and high-pressure (orthorhombic) phases was obtained by fitting the $P$-$V$ data to a third-order Vinet equation of state Vinet et al. (1987), as shown in Figure 3. The fits provided a bulk modulus value of $B_{0}$ = 52(7) GPa with a derivative $B_{0}^{\prime}$ = 9.2 for the cubic phase and $B_{0}$ = 68(1) GPa with $B_{0}^{\prime}$ = 4.8 for the orthorhombic phase. The compressibility of Ag3AuTe2 is comparable to other reported Ag ternary chalcogenides such as AgSbTe2 and AgBiSe2.Berri et al. (2013); Ko et al. (2014); Elafy et al. (2019) The high strain tolerance of Ag3AuTe2 bestow prospects for further investigation into the high-pressure band gap dynamics in light of the previous computational findings which suggested the possibility of strain-induced band engineering. Complementary techniques such as in situ high-pressure UV-Vis and resistivity measurements would provide valuable insights into its topological nature.
III.3 Electronic properties
Figure 5: Electrical resistivity of Ag3AuTe2 is plotted against temperature. The inset Arrhenius plot was obtained by applying a linear regression model with the least-squares method, which returned an activation energy of approximately 0.20 eV.
Electronic properties of Ag3AuTe2 were examined experimentally by transport measurements. Figure 5 shows the electrical resistivity of Ag3AuTe2 as a function of temperature with the corresponding inset Arrhenius plot generated using the activation law. The material exhibits a semiconductor behavior and high resistivity, nearing the megaohm range at around 100 K. The magnitude of the activation energy $E_{a}$ aligns well with the previously proposed DFT-calculated band gaps.Faizan et al. (2016); Sánchez-Martínez et al. (2019) Young et al.Young et al. (2000) reported similar experimental results, albeit with a slightly larger $E_{a}$ of 0.3 eV and no evidence for sample purity. Differences in the estimated $E_{a}$ are within the typical error range for resistivity measurements stemming from factors such as imperfections in sample geometry, inhomogeneity, and possible in-gap defects. Additionally, a resistivity kink near 275 K in Figure 5, also observed by Young et al.Young et al. (2000) with no explanation, is likely an artifact from the freezing of adsorbed water.
Transport measurements confirm that Ag3AuTe2 shows semiconductor-like temperature dependence and a band gap of at least 0.2 eV, accounting for potential defect hopping effects. To evaluate petzite’s potential for strain-induced band crossing, accurate determination as well as prediction of band gap is important. Subsequent DFT analysis explores the band structure and possible band gap modulation through lattice parameter adjustments.
Figure 6: DFT-calculated band structures with spin-orbit coupling of Ag3AuTe2 at ambient pressure and the maximum experimental compression before dissociation ($a=9.94$, 13 GPa). A semi-metallic behavior is predicted by the PBE functional, while the more accurate band gap predicting method, mBJ, predicts a gap even under pressure.
% 0 1 2 3 4
PBE 0.093 0.049 0.013 0 0
mBJ 0.603 0.516 0.438 0.383 0.272
Table 1: DFT-calculated band gaps of Ag3AuTe2 with an incremental reduction in the lattice parameter. Band gaps are in the unit of eV. The PBE and mBJ gaps decrease with reduced lattice parameters, but at 4% compressive strain, the PBE gap achieves metallicity while the mBJ gap remains open.
Electronic properties were computationally investigated by density functional theory (DFT) band structure calculations. Figure 6 illustrates the calculated band structure of Ag3AuTe2 with a reduction of lattice constant up to 4% compressive strain, using the Perdew-Burke-Ernzerhof parameterization (PBE) and modified Becke-Johnson (mBJ), including spin-orbit coupling (SOC). Calculations without SOC are provided in the supplementary material. SOC breaks Kramer’s degeneracy and splits the band structure for different spin components, but it does not induce qualitative differences in the band dispersion and the gap. Table 1 summarizes the results, indicating that at ambient pressure ($a=10.38$), the indirect band gap between the valence band near $N$ and the conduction band around $\Gamma$ is 0.093 and 0.603 eV for PBE and mBJ functionals, respectively. mBJ calculations predict that negative strain (compression) decreases the band gap while conserving the overall shape and features of the band structure, and the material remains to be an insulator at the maximal compressive strain (-4%). The PBE functional, known to underestimate band gaps, also predicts decreasing band gaps with negative strain. However, the conduction and valence band overlap near $\Gamma$ along the $\Gamma-H$ cut before reaching -4% strain, transitioning the material into a metal under pressure. The SOC splitting strengthens with increased strain and the predicted SOC effect by mBJ functional is less significant than that using PBE method. Our PBE results are mostly consistent with previous DFT studiesSánchez-Martínez et al. (2019); Gosálbez-Martínez et al. (2023). However, we expect the mBJ functional to provide more accurate calculation of the band structure in general, and the shape and splitting of the conduction band around $\Gamma$, in particular, are qualitatively different from the PBE predictions.
To unambiguously visualize the atomic contribution near the band edges, Ag, Te, and Au are color-coded by magenta, cyan, and yellow, respectively, in Figure 6 for PBE and mBJ functionals on strained and ambient lattices. The band structures are highly hybridized. The valence bands are primarily composed of an equal mixture of Ag-d and Te-p orbitals, indicated by blue shades, and a smaller contribution from Ag-s orbitals. The conduction bands largely consist of Au-s and d orbitals, although contributions from other atoms remain substantial. In general, the band character of the conduction bands becomes mixed with the valence bands as the band gap is reduced.
Ag3AuTe2, along with the isostructural Ag3AuSe2, were theoretically analyzed in Refs. Sánchez-Martínez et al., 2019 and Won et al., 2022. In particular, Ref. Sánchez-Martínez et al., 2019 used DFT calculations to predict that both compounds feature a gap that narrows as a function of applied pressure. The low-energy bands under pressure are centered at the $\Gamma$ point and involve fourfold degenerate chiral multifold fermions with nontrivial Chern numbers. Ref. Sánchez-Martínez et al., 2019 developed a low-energy $\mathbf{k}\cdot\mathbf{p}$ approximation of the bands near $\Gamma$ and showed that in Ag3AuSe2 the optical conductivity under pressure was dominated by contributions from the multifold fermions Bradlyn et al. (2016); Flicker et al. (2018). Due to the qualitative similarity between the electronic bands of Ag3AuSe2 and Ag3AuTe2, we expect that optical properties of Ag3AuTe2 are also due to the multifold fermions.
III.4 Optical Properties
Figure 7: Tauc plot of Ag3AuTe2 was obtained from the diffuse reflectance measurement. The optical band gap energy was estimated by the intersection of the baseline (blue) and fitted line (red) since the baseline was not flat to zero. The approximate band gap is 0.58 eV.
Optical characterization of Ag3AuTe2 was performed by UV-Vis diffuse reflectance measurements at room temperature. The obtained reflectance spectra were then transformed into absorption spectra using the Kubelka-Munk function,Kubelka and Munk (1931) expressed as
$F(R_{\infty})=\frac{(1-R_{\infty})^{2}}{2R_{\infty}},$ (1)
where $R_{\infty}$ represents the reflectance of a medium with infinite thickness. Replacing the absorption coefficient $\alpha$ in the Tauc equationTauc et al. (1966) with $R_{\infty}$ yields the expression (2)
$(F(R_{\infty})\cdot\\ h\nu)^{1/\gamma}=B(h\nu-E_{g}),$ (2)
where the $\gamma$ factor is equal to 1/2 for a direct gap and 2 for an indirect gap.
Using the above relations with $\gamma=1/2$ resulted in a Tauc plot with a gentle, rather than a steep rise at approximately 0.6 eV, as shown in Figure 7, consistent with the predicted mBJ ambient pressure indirect gap. However, the PBE gap falls short of the optical band gap by orders of magnitude, as expected due to the tendency of PBE functional to underestimate band gaps. The Arrhenius activation energy $E_{a}$ (0.2 eV) is lower than but comparable to the optical gap in magnitude. Discrepancies in the experimental gap values may arise from factors such as errors in sample dimensions for transport measurement and a rough approximation of the optical absorption onset. The correspondence between the optical and mbJ gap suggests promising avenues for exploiting pressure-induced band modulation of Ag3AuTe2.
IV Conclusions
We investigated the synthesis, electronic and optical properties, as well as the compressibility of polycrystalline Ag3AuTe2. High-pressure XRD measurements revealed a pressure-induced structural transition of petzite around 16 GPa, shifting from a cubic to an orthorhombic Ag2Te-type structure. Transport results corroborated Young et al.’s observations,Young et al. (2000) highlighting the semiconductor nature of the telluride with an Arrhenius activation energy of 0.2 eV, which falls within the predicted gap range by PBE and mBJ DFT calculations. The absorption band edge determined by diffuse reflectance measurements was in excellent agreement with the mBJ gap. First-principles calculations using PBE and mBJ functionals predicted a gap reduction with decreasing lattice parameters, reaching zero and 0.3 eV gap at -4% strain, respectively. Our high-pressure XRD experiments demonstrated a lattice parameter reduction of over 4% in Ag3AuTe2, suggesting a viable band engineering avenue for Ag3AuTe2 towards zero gap under applied pressure.
V Acknowledgments
Crystal growth, transport, and optical characterization were supported by the Center for Quantum Sensing and Quantum Materials, an Energy Frontier Research Center funded by the U. S. Department of Energy (DOE), Office of Science (SC), Basic Energy Sciences (BES) under Award DE-SC0021238. First-principles calculations conducted by RZ and CP were supported by the DOE-SC-BES under Award DE-AC02-76SF00515. The authors acknowledge using facilities at the Materials Research Laboratory Central Research Facilities, University of Illinois, partially supported by NSF through the University of Illinois Materials Research Science and Engineering Center (DMR-1720633). High-pressure X-ray diffraction measurements were performed at HPCAT (Sector 16), Advanced Photon Source (APS), Argonne National Laboratory. HPCAT operations are supported by the DOE-National Nuclear Security Administration (NNSA) Office of Experimental Sciences. The beamtime and contributions of RK and RJH were made possible by the Chicago/DOE Alliance Center (CDAC), which is supported by DOE-NNSA (Grant No. DE-NA0003975). RK and RH were also supported by NSF DMR-2119308 and DMR-2104881. COMPRES supported use of the gas loading system under NSF Cooperative Agreement No. EAR-1606856 and by GSECARS through NSF Grant No. EAR-1634415 and DOE Grant No. DE-FG02-94ER14466. The Advanced Photon Source is a DOE Office of Science User Facility operated for the DOE Office of Science by Argonne National Laboratory under Contract No. DE-AC02-06CH11357. This research used the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award BES-ERCAP0027203.